# VICtoR: Learning Hierarchical Vision-Instruction Correlation Rewards for Long-horizon Manipulation

### v1

@@ Abstract
We study reward models for long-horizon manipulation tasks by learning from action-free videos and language instructions, which we term the visual-instruction correlation (VIC) problem. Recent advancements in cross-modality modeling have highlighted the potential of reward modeling through visual and language correlations. However, existing VIC methods face challenges in learning rewards for long-horizon tasks due to their lack of sub-stage awareness, difficulty in modeling task complexities, and inadequate object state estimation. To address these challenges, we introduce VICtoR, a novel hierarchical VIC reward model capable of providing effective reward signals for long-horizon manipulation tasks. VICtoR precisely assesses task progress at various levels through a novel stage detector and motion progress evaluator, offering insightful guidance for agents learning the task effectively. To validate the effectiveness of VICtoR, we conducted extensive experiments in both simulated and real-world environments. The results suggest that VICtoR outperformed the best existing VIC methods, achieving a 43% improvement in success rates for long-horizon tasks.
@@ 1 Introduction
Reinforcement learning (RL) has been extensively studied for long-horizon manipulation [1, 2, 3]. However, crafting reward functions for these tasks is complex, as it typically requires access to true states or significant domain expertise. Consequently, there is an urgent need for a robust and accurate reward model for these tasks. Previous research has explored modeling reward functions via robotic expert demonstrations [4, 5], goal-images [6, 7], and human demonstrations [8, 7, 9]. Unfortunately, the required task specification materials, such as goal-images, remain costly and impractical.
Recently, methods that consider the vision-instruction correlation (VIC) as reward signals have emerged, providing a more accessible way for task specification through language. Specifically, these VIC methods view the reward modeling as a regression or classification problem and train the reward model on the given action-free demos and instructions. Pioneer approaches [10, 11, 12, 13, 14] studied using pre-trained Vision Language Models (VLMs) or Video Language Models, like CLIP [15], to generate rewards by assessing the similarity between visual observations and language goals. Ma et al. [16] explore the possibility of using human video and language data to pre-train a reward model, later fine-tuned with in-domain data. However, existing VIC studies are all limited to short-horizon tasks, such as “picking up a block" or “opening a drawer".
@@
To this end, we propose VICtoR, a hierarchical VIC reward model for long-horizon manipulation. Figures 1 and 2 depict its concept and architecture, respectively. VICtoR learns effective rewards for long-horizon tasks by hierarchically assessing overall progress. Specifically, it decomposes long-horizon tasks (high-level) into stages (mid-level) and motions and progress (low-level, robot primitives), determining the reward by considering task progress at different granularities. By treating each long-horizon task as an arbitrary permutation of available motions, VICtoR, trained only on action-free motion videos, can be easily adapted to unseen long-horizon tasks.
To achieve this, VICtoR first queries GPT-4 to break down a long-horizon task into stages to acquire essential task knowledge, including the expected object statuses and required motions for completion at each stage. Then, a stage detector (a local-scale VLM) retrieves the current object statuses from the current observation and verifies if they match the condition of a specific stage in the task knowledge (Challenge #$\#1$). Next, from a list of required motions to complete the stage, a motion progress evaluator (a whole-scale VLM) determines the current motion and assesses the in-motion progress (Challenges #$\#2$ and #$\#3$) by evaluating the correlations between visual observations and each motion instruction. As a result, VICtoR can generate informative rewards to complete long-horizon tasks.
We conducted extensive experiments in both simulated and real-world settings. In the simulation, we designed ten tasks of varying horizons, including those where each step’s success depends on the success of the previous stage. For real-world experiments, we used the dataset collected from XSkill[17]. Compared to the best prior VIC reward model, VICtoR achieved a 25% improvement in the average success rate across all tasks and a 43% improvement in the more challenging tasks (top half of tasks considering motion’s number). Additionally, our detailed ablation studies emphasized the importance of our hierarchical architecture and training objectives. Visualizations of the learned embedding space and rewards demonstrated that VICtoR effectively assesses task progress.
@@ 2 Related Work
With great ability of reasoning, Large Language Models (LLMs) have been utilized for robotics in navigation [18], task planning [19, 20], code-completion for policies [21, 22], and manipulation [23, 24, 25, 26, 27]. These works leverage LLMs’ ability to translate high-level instructions into actionable sequences and break them down into sub-orders for precise agent control. Additionally, recent advancements have leveraged Vision-Language Models (VLMs) to enhance environmental understanding for agents. With a profound understanding of the physical world, VLMs facilitate the works in navigation, robotic control, and monitoring [28, 29, 30, 31, 32]. Unlike previous methods, for reward modeling, VICtoR uses LLMs to break down tasks and VLMs to assess visual observation with motion descriptions, allowing our method to better assess the task’s progress and distinguish it with different granularity.
@@
Long-horizon manipulation is a long-standing problem. Traditional Hierarchical Reinforcement Learning (HRL) methods [33, 34, 2, 35, 36] address this challenge by decomposing complex tasks into hierarchical levels or sequential orders, where higher-level policies make decisions that guide the selection of parameters for lower-level policies or pre-trained actions. Task and Motion Planning (TAMP) methods [37, 38, 39] attempt to assemble and schedule well-trained motions to accomplish long-horizon tasks. Additionally, Long-horizon Imitation Learning [40] has been proposed to learn long-horizon tasks from expert demonstrations. Unlike these approaches, VICtoR aims to learn the reward model for long-horizon tasks from action-free demonstrations and task instructions. Our reward model is designed to be independent of policy design and can be applied to RL algorithms to learn long-horizon manipulation tasks.
@@ 3 Preliminaries
In the context of long-horizon manipulation, we define motion as as the primitive movement that a robot can make, such as reaching or grasping an object. In addition, stage refers to a high-level interaction with an object in the environment, requiring the consideration of complex objects’ status, and can encompass multiple motions. For instance, the stage "open the drawer" involves the motions "reach the drawer handle" and "pull out the drawer." A comprehensive list of motions and stages is detailed in Appendix B.1.
@@ 4 Method
Figure 2: Training and inference pipeline of VICtoR. VICtoR is trained using motion-level videos with language annotations and object state labels. It first decomposes the task into task knowledge for decomposed stages, conditional object states, and motions. Then, it uses Stage Detector to identify the stage, and a Motion Progress Evaluator (VLM) to detect the motion and in-motion progress
@@ 4.1 Task Knowledge Generation
To make long-horizon tasks more achievable and manageable, we leverage the reasoning ability in Large Language Models (LLMs) to decompose and analyze the tasks. Previous research [19, 20] introduced LLMs into policy learning by planning sub-goals or actions for explicit guidance to enhance long-horizon capabilities. Building upon this foundation, we explore more nuanced usage. Our goal is to decompose a given training task into distinct sub-stages, define the specific conditions or states of each object at these stages, and identify the motions required by the agent to advance to the next stage. To achieve this, we employ GPT-4 to analyze and construct task knowledge. Figure 2 illustrates an example of task knowledge generation for the task “move the block into the closed drawer.” Details of our prompt design and implementation are provided in Appendix F, and additional experiments for robustness are described in Appendix A.2.
@@ 4.2 Stage Detection
To detect object states, our Stage Detector first leverages a language-conditioned object detector $D$, such as MDETR [41], to identify objects in the image observation $o$. Essentially, we crop out the bounding box $D(o,c)$ generated from the detector $D$ based on the current observation $o$ and the queried object $c$. Subsequently, these cropped images are fed into our object state classifier $P$. The classifier predicts the likelihood of categorical object state for each cropped image as $P(L^{c}|D(o,c))$, where $L^{c}$ represents the language description of possible object states for object $c$. We summarize the training details in Appendix B.2.
@@
After extracting all the object states in the environment, our Stage Detector can then determine the current stage by comparing these states with the task knowledge. A stage is identified only if, and only if, all object states align with the task knowledge for a specific stage. For example, as illustrated in Figure 2, if the drawer is currently closed, we compare this condition with the states specified in each stage of the task knowledge. Upon finding that the drawer is expected to be closed only at the first stage, we conclude that the agent is at the first stage. This comparison should be applied to every conditioned object for the tasks. If the extracted set of object states does not match any stage, the agent will be considered to be in the initial stage.
@@ 4.3 Motion Progress Evaluator
Since a stage may require multiple motions to complete, relying merely on the stage reward is insufficient for learning complex tasks. To this end, we introduce the Motion Progress Evaluator (MPE) to provide the reward signals within the stage. As discussed in Section 4.1, each detected stage comes with a list of motion descriptions in the task knowledge. Building on this, the MPE assesses the agent’s current motion and progress within the motion, taking into account these descriptions. This MPE model yields a more detailed reward model that accurately reflects progress within stages.
@@
To effectively evaluate progress within a stage, we expect three key capabilities in the MPE model: (1) the ability to discern the temporal difference of frames extracted from the same video, (2) the capacity to identify the motion in which the agent is engaged, and (3) the capability to assess the progress of motion completion. To this end, we introduce objectives based on InfoNCE [42] to guide our MPE model, focusing on time contrast, motion contrast, and language-frame contrast.
@@
To capture features relevant to physical interaction and sequential decision-making, we apply the time contrastive loss $\mathcal{L}_{tcn}$ to the base embeddings, as inspired by [43, 44]. Essentially, images that are temporally closer should have more similar representations (embeddings) than those that are temporally distant or from different videos. The loss $\mathcal{L}_{tcn}$ can be formulated by
@@ 5.1 Experiment Settings
Most existing manipulation benchmarks [46, 47, 48] do not support long-horizon tasks where the success of each step depends on the outcomes of previous steps. While Calvin [49] addresses the aforementioned issue, the tasks and action spaces (e.g., rotate) are too challenging for VIC reward models trained on action-free videos. Consequently, we developed a new environment to test different VIC reward models for long-horizon manipulation. For policy training in this environment, we use a ground-truth object detector to speed up the process, as we need to run multiple tasks with multiple seeds. For real-world experiments, we train and test reward models using the XSkill [17] dataset. Details on the environment, tasks, and data are in Appendix B.1 and Figure 3.
@@ 5.2 Policy Learning with VIC Rewards
In this section, we train the Proximal Policy Optimization algorithm [51] using various reward functions. We utilize a sparse reward as a signal for task success and shape the reward function using different reward models. Our experiments are divided into two sets: initially, we train on 3 single-stage tasks with varying numbers of motions, and subsequently, on 7 tasks with varying numbers of stages. As shown in Tables 1 and 2, VICtoR consistently outperforms current baselines, achieving a 25% average performance gain. This advantage increases to 43% in harder tasks with more than three motions, highlighting VICtoR’s effectiveness in training for long-horizon tasks. The significant benefits of the motion design are also evident when compared to the task-level baseline, especially in more complex tasks. Furthermore, as Table 1 demonstrates, even at the task level, VICtoR outperforms other baselines, showcasing the effectiveness of our training objective detailed in Section 4.3. Training details and curves are provided in Appendix B.5, and the qualitative analyses for challenges outlined in Figure 1 are also provided in Appendix A.1.
@@
Table 2: Experiment of tasks with different numbers of stages and motions. Extending from 1, this table includes seven additional tasks with varying numbers of stages and motions. It shows that VICtoR outperforms prior methods, and the differences become larger as the complexity of the tasks increases, demonstrating the effectiveness of our design of stage and motion rewards. Results are averaged over three different seeds.
@@ 5.4 Real World Experiments
To demonstrate VICtoR’s practical efficacy, we visualized the potential curves $\phi(o_{t})$ for two scenarios: one with videos that match the corresponding instructions, and another with incorrect videos using the same instructions. As illustrated in Figure 4, for the correct actions, the potential curves show that VICtoR effectively identifies task progress in real-world scenarios by increasing the potential as the agent completes tasks, a capability not matched by previous VIC reward models. For the incorrect actions, as the agent moves from the right to the left side to close the drawer, VICtoR‘s visualization shows an initial increase in potential. This increase is logical as these movements approach the light, aligning with the first stage of the instruction, “open the light.” However, as the agent continues toward the drawer, VICtoR recognizes the incorrect task and begins to decrease the reward, effectively deterring the agent’s movement. This highlights VICtoR’s ability to analyze agent movement and task progress accurately. Additional visualizations for various tasks are provided in Appendix E.
@@ 6 Conclusion
Appendix A Additional Experiments
A.1 Qualitative Analysis
To further analyze different reward models, we also examine the potential curves for various reward models to substantiate our statement shown in Figure 1. In the qualitative analysis, we aim to demonstrate our statement regarding (1) confusion arising from varying task difficulties, and (2) the lack of explicit object state estimation.
For statement (1), we plot out the potential curves from 3 different reward models on 4 tasks within variant task complexities. The result in Figure 6 reveals that VICtoR is capable of generating effective reward curves on composed and complex tasks as other baseline models fail to provide progressive signals for agents when the task horizon grows.
To demonstrate statement (2), we visualize the reward by giving text goals for long-horizon tasks (open light then open drawer then reach blue block) on misordered action (reach blue block) videos to see the difference in reward models. The result in Figure 7 indicates that VICtoR is capable of determining the current state of environmental objects to decide the level of potential while LIV simply generates progressive signals on each task.

### v2

@@ Abstract
We study reward models for long-horizon manipulation by learning from action-free videos and language instructions, which we term the visual-instruction correlation (VIC) problem. Existing VIC methods face challenges in learning rewards for long-horizon tasks due to their lack of sub-stage awareness, difficulty in modeling task complexities, and inadequate object state estimation. To address these challenges, we introduce VICtoR, a novel hierarchical VIC reward model capable of providing effective reward signals for long-horizon manipulation tasks. Trained solely on primitive motion demonstrations, VICtoR effectively provides precise reward signals for long-horizon tasks by assessing task progress at various stages using a novel stage detector and motion progress evaluator. We conducted extensive experiments in both simulated and real-world datasets. The results suggest that VICtoR outperformed the best existing methods, achieving a 43% improvement in success rates for long-horizon tasks.
@@ 1 Introduction
Reinforcement learning (RL) has been extensively studied for long-horizon manipulation [1; 2; 3]. However, crafting reward functions for these tasks is complex, as it typically requires access to true states or significant domain expertise. Consequently, there is an urgent need for a robust and accurate reward model for these tasks. Previous research has explored modeling reward functions via robotic expert demonstrations [4; 5], goal-images [6; 7], and human demonstrations [8; 7; 9]. Unfortunately, the required task specification materials, such as goal-images, remain costly and impractical.
Recently, methods that consider the vision-instruction correlation (VIC) as reward signals have emerged, providing a more accessible way for task specification through language. Specifically, these VIC methods view the reward modeling as a regression or classification problem and train the reward model on the given action-free demos and instructions. Pioneer approaches [10; 11; 12; 13; 14] studied using pre-trained Vision Language Models (VLMs) or Video Language Models, like CLIP [15], to generate rewards by assessing the similarity between visual observations and language goals. Additionally, Ma et al. [16] explore the possibility of using human video and language data to pre-train a reward model, later fine-tuned with in-domain data. However, regardless of whether diverse data is used for pre-training, existing VIC studies still require full-length data to retrain the model. Notably, these methods are constrained to short-horizon tasks, such as “picking up a block".
@@
We propose VICtoR, a hierarchical VIC reward model for long-horizon manipulation. Figures 1 and 2 depict its concept and architecture. VICtoR learns effective rewards for long-horizon tasks by hierarchically assessing overall progress. Specifically, it decomposes long-horizon tasks (high-level) into stages (mid-level) and motions and progress (low-level, robot primitives), determining the reward by considering task progress at different granularities. Meanwhile, the hierarchical feature sets VICtoR free from the requirements of full-length demonstrations and further grants better generalizability. By treating each long-horizon task as an arbitrary permutation of available motions, VICtoR, trained only on action-free motion videos, can be adapted to unseen long-horizon tasks.
To minimize human effort, we leverage GPT-4’s semantic understanding to extract task knowledge (expected object statuses and required motions). Then, a stage detector retrieves the current object statuses from the observation and verifies if they match the condition of a specific stage in the task knowledge (Challenge #$\#1$). Next, from a list of required motions to complete the stage, a motion progress evaluator determines the current motion and assesses the in-motion progress (Challenges #$\#2$ and #$\#3$) by evaluating the correlations between visual observations and each motion instruction. Consequently, VICtoR can generate informative rewards to complete long-horizon tasks.
We conducted extensive experiments in both simulated and real-world settings. In the simulation, we designed ten tasks of varying horizons, including those where each step’s success depends on the success of the previous stage. For the real-world experiments, we used the dataset collected from XSkill [17] and visualized the trained reward curves in comparison to other baselines. Compared to the best prior VIC reward model, VICtoR achieved a 25% improvement in the average success rate across all tasks and a 43% improvement in the more challenging tasks (tasks which consisted of more than 4 motions). Additionally, our detailed ablation studies emphasized the importance of our hierarchical architecture and training objectives. Visualizations of the learned embedding space and rewards demonstrated that VICtoR effectively assesses task progress.
@@ 2 Related Work
With great ability of reasoning, Large Language Models (LLMs) have been utilized for robotics in navigation [18], task planning [19; 20], code-completion for policies [21; 22], and manipulation [23; 24; 25; 26; 27]. These works leverage LLMs’ ability to translate high-level instructions into actionable sequences and break them down into sub-orders for precise agent control. Additionally, recent advancements have leveraged Vision-Language Models (VLMs) to enhance environmental understanding for agents. With a profound understanding of the physical world, VLMs facilitate the works in navigation, robotic control, and monitoring [28; 29; 30; 31; 32]. Unlike previous methods, for reward modeling, VICtoR uses LLMs to break down tasks and VLMs to assess visual observation with motion descriptions, allowing our method to better assess the task’s progress and distinguish it with different granularity.
@@
Long-horizon manipulation is a long-standing problem. Traditional Hierarchical Reinforcement Learning (HRL) methods [33; 34; 2; 35; 36] address this challenge by decomposing complex tasks into hierarchical levels or sequential orders, where higher-level policies make decisions that guide the selection of parameters for lower-level policies or pre-trained actions. Task and Motion Planning (TAMP) methods [37; 38; 39] attempt to assemble and schedule well-trained motions to accomplish long-horizon tasks. Additionally, Long-horizon Imitation Learning [40; 41] has been proposed to learn long-horizon tasks from expert demonstrations. Unlike these approaches, VICtoR aims to learn the reward model for long-horizon tasks from action-free demonstrations and task instructions. Our reward model is designed to be independent of policy design and can be applied to RL algorithms to learn long-horizon manipulation tasks.
@@ 3 Preliminaries
In the context of long-horizon manipulation, we define motion as the primitive movement that a robot can make, such as reaching or grasping an object. In addition, stage refers to a high-level interaction with an object in the environment, requiring the consideration of complex objects’ status, and can encompass multiple motions. For instance, the stage "open the drawer" involves the motions "reach the drawer handle" and "pull out the drawer." A comprehensive list of motions and stages is detailed in Appendix C.1.
@@ 4 Method
Figure 2: Training and inference pipeline of VICtoR. VICtoR is trained using motion-level videos with language annotations and object state labels. It first decomposes the task into task knowledge for decomposed stages, conditional object states, and motions. Then, it uses Stage Detector to identify the stage, and a Motion Progress Evaluator (VLM) to detect the motion and in-motion progress.
@@ 4.1 Task Knowledge Generation
To make long-horizon tasks more achievable and manageable, we leverage the reasoning ability in Large Language Models (LLMs) to decompose and analyze the tasks. Previous research [19; 20] introduced LLMs into policy learning by planning sub-goals or actions for explicit guidance to enhance long-horizon capabilities. Building upon this foundation, we explore more nuanced usage. Our goal is to decompose a given training task into distinct sub-stages, define the specific conditions or states of each object at these stages, and identify the motions required by the agent to advance to the next stage. To achieve this, we employ GPT-4 to analyze and construct task knowledge. Figure 2 illustrates an example of task knowledge generation for the task “move the block into the closed drawer.” Details of our prompt design and implementation are provided in Appendix G, and additional experiments for robustness are described in Appendix B.2.
@@ 4.2 Stage Detection
To detect object states, our Stage Detector first leverages a language-conditioned object detector $D$, such as MDETR [42], to identify objects in the image observation $o$. Essentially, we crop out the bounding box $D(o,c)$ generated from the detector $D$ based on the current observation $o$ and the queried object $c$. Subsequently, these cropped images are fed into our object state classifier $P$. The classifier predicts the likelihood of categorical object state for each cropped image as $P(L^{c}|D(o,c))$, where $L^{c}$ represents the language description of possible object states for object $c$. We summarize the training details in Appendix C.2.
@@
After extracting all the object states in the environment, our Stage Detector can then determine the current stage by comparing these states with the task knowledge. A stage is identified if, and only if, all object states align with the task knowledge for a specific stage. For example, as illustrated in Figure 2, if the drawer is currently closed, we compare this condition with the states specified in each stage of the task knowledge. Upon finding that the drawer is expected to be closed only at the first stage, we conclude that the agent is at the first stage. This comparison should be applied to every conditioned object for the tasks. If the extracted set of object states does not match any stage, the agent will be considered to be in the initial stage.
@@ 4.3 Motion Progress Evaluator
Since a stage may require multiple motions to complete, relying merely on the stage reward is insufficient for learning complex tasks. To this end, we introduce the Motion Progress Evaluator (MPE) to provide the reward signals within the stage. As discussed in Section 4.1, each detected stage comes with a list of motion descriptions in the task knowledge. Building on this, the MPE assesses the agent’s current motion and progress within the motion, taking into account these descriptions. This MPE model yields a more detailed reward signal that accurately reflects progress within stages.
@@
To effectively evaluate progress within a stage, we expect three key capabilities in the MPE model: (1) the ability to discern the temporal difference of frames extracted from the same video, (2) the capacity to identify the motion in which the agent is engaged, and (3) the capability to assess the progress of motion completion. To this end, we introduce objectives based on InfoNCE [43] to guide our MPE model, focusing on time contrast, motion contrast, and language-frame contrast. The functions of the three objectives are illustrated in Figure 7 in Appendix.
@@
To capture features relevant to physical interaction and sequential decision-making, we apply the time contrastive loss $\mathcal{L}_{tcn}$ to the base embeddings, as inspired by [44; 45]. Essentially, images that are temporally closer should have more similar representations (embeddings) than those that are temporally distant or from different videos. The loss $\mathcal{L}_{tcn}$ can be formulated by
@@ 5.1 Experiment Settings
Most existing manipulation benchmarks [47; 48; 49] do not support long-horizon tasks where the success of each step depends on the outcomes of previous steps. While Calvin [50] addresses the aforementioned issue, the tasks and action spaces (e.g., rotate) are too challenging for VIC reward models trained on action-free videos. Consequently, we developed a new environment to test different VIC reward models for long-horizon manipulation. For policy training in this environment, we use a ground-truth object detector to speed up the process, as we need to run multiple tasks with multiple seeds. For real-world experiments, we train and test reward models using the XSkill [17] dataset. Details on the environment, tasks, and data are in Appendix C.1 and Figure 3.
@@ 5.2 Policy Learning with VIC Rewards
In this section, we train the Proximal Policy Optimization algorithm [52] using various reward functions on the simulated environment. We utilize a sparse reward as a signal for task success and shape the reward function using different reward models. Our experiments are divided into two sets: initially, we train on 3 single-stage tasks with varying numbers of motions, and subsequently, on 7 tasks with varying numbers of stages. As shown in Tables 1 and 2, VICtoR consistently outperforms current baselines, achieving a 25% average performance gain. This advantage increases to 43% in harder tasks with more than three motions, highlighting VICtoR’s effectiveness in training for long-horizon tasks. The significant benefits of the motion design are also evident when compared to the task-level baseline, especially in more complex tasks. Furthermore, as Table 1 demonstrates, even at the task level, VICtoR outperforms other baselines, showcasing the effectiveness of our training objective detailed in Section 4.3. Training details and curves are provided in Appendix C.4.
@@
Table 2: Experiment of tasks with different numbers of stages and motions. Extending from Table 1, this table includes seven additional tasks with varying numbers of stages and motions. It shows that VICtoR outperforms prior methods, and the differences become larger as the complexity of the tasks increases, demonstrating the effectiveness of our design of stage and motion rewards. Results are averaged over three different seeds.
@@ 5.3 Ablation Studies
5.4 Qualitative Analysis
We plot the potential curves from three different reward models across four tasks with varying complexities (with longer horizons on the right). The results in Figure 4 demonstrate that VICtoR effectively generates reward curves for complex, composed tasks, while other baseline models fail to provide progressive signals as task horizons grow. Notably, while prior reward models were trained on full-length, long-horizon demonstrations, VICtoR was only trained on motion-level data yet still achieved the most optimal results. We also put the analysis for misordered action in Appendix B.1.
@@ 5.5 Real World Experiments
To demonstrate VICtoR’s practical efficacy, we visualized the potential curves $\phi(o_{t})$ for two scenarios: one with videos that match the corresponding instructions, and another with incorrect videos using the same instructions. As illustrated in Figure 5, for the correct actions, the potential curves show that VICtoR effectively identifies task progress in real-world scenarios by increasing the potential as the agent completes tasks, a capability not matched by previous VIC reward models. For the incorrect actions, as the agent moves from the right to the left side to close the drawer, VICtoR‘s visualization shows an initial increase in potential. This increase is logical as these movements approach the light, aligning with the first stage of the instruction, “open the light.” However, as the agent continues toward the drawer, VICtoR recognizes the incorrect task and begins to decrease the reward, effectively deterring the agent’s movement. This highlights VICtoR’s ability to analyze agent movement and task progress accurately. Additional visualizations for various tasks are provided in Appendix F.
@@ 6 Conclusion
Acknowledgements
This work was supported in part by the National Science and Technology Council, Taiwan, under Grant NSTC 113-2634-F-002-007 and NSTC 113-2813-C-002 -031 -E. We also thank all reviewers and area chairs for their valuable comments and positive recognition of our work during the review process.
@@ Acknowledgements
Appendix A Illustrations for Loss Objectives
Figure 7: In-motion and cross-motion Contrastive Objectives: We illustrate how the three contrastive loss functions we applied manipulate the distances and relationships between instruction representations and encoded visual observations within the embedding space.
Appendix B Additional Experiments
B.1 Qualitative Analysis - Misordered Action
We visualize the reward by giving text goals for long-horizon tasks (open light then open drawer then reach blue block) on misordered action (reach blue block) videos to see the difference in reward models. The result in Figure 8 indicates that VICtoR is capable of determining the current state of environmental objects to decide the level of potential while LIV simply generates progressive signals on each task.